# Computational Modelling of Thixotropic Multiphase Fluids

### v1

@@ 3.1 Validation of the numerical methodology
We conduct a series of benchmark simulations to ensure the accuracy and robustness of our implementation. First addressing the correct description of the surface tension and contact angles in static and dynamic cases. Then, we validate the thixotropic model by simulating a transient viscosity case. The results of these simulations are compared with theoretical predictions and previous numerical results
3.1.1 Static validation: Surface tension and contact angles
@@
As several research studies have shown (see [22, 23, 24, 15, 25, 17]), within the validation of numerical methods for multiphase flows, the analysis of the surface tension play an important role. The Young Laplace equation indicates that the difference in pressure between the droplet ($P_{d}$) and the external fluid ($P_{f}$) is directly proportional to the surface tension ($\sigma$) and inversely proportional to the radius ($R$), i.e., $P_{d}-P_{f}=2\sigma/R$. We model the effect of surface tension starting with a square-shaped particle array, which evolves in time to a circular droplet. The size of the initial square varies to produce a range of equilibrium radius. For these simulations, we use $\sigma=0.25$ and $\sigma=1$, with an equilibrium particle density $n_{eq}=25$ and $k_{b}T\to 0$. In Figure 1(a), we present the evolution of the pressure difference as a function of curvature ($1/R$). Consistent with the Young-Laplace equation, the slopes obtained from the linear fitting, $0.53$ and $2.08$ corresponds to $\sim 2\sigma$. We further validate the consistency of our method, measuring the deformation of an initially elliptical droplet as it retracts to a circular shape due to interfacial tension effects[30, 31, 32, 33]. The initial ellipse is defined by the semi-major axis $A$ and the semi-minor axis $B$, with an initial Taylor deformation parameter[29] ${\mathcal{D}}_{t=0}=(A-B)/(A+B)$. The retraction process is governed by the balance between surface tension and viscous forces. The evolution of the of the Taylor parameter $\mathcal{D}$ can be described by[30, 31, 32]
$\mathcal{D}=\mathcal{D}_{t=0}\exp\left(-\frac{\sigma}{\eta R_{0}}\frac{40(\lambda+1)}{(2\lambda+3)(19\lambda+16)}t\right),$ (15)
@@
where $R_{0}$ the final circle radius, $\lambda$ the ratio between the viscosity of the phases and $t$ the stretching time. Here, we use three different viscosity ratios $\lambda=[1,5,10]$, and use Principal Component Analysis (PCA) to compute the eigenvalues and eigenvectors of the droplets particles, providing the magnitude and orientation of the ellipse’s principal axes. In Figure 1(b), we compare our results for the time evolution of $\mathcal{D}$ with (15), evidencing a good agreement between the numerical and theoretical results.
@@
We now validate the model’s ability to consistently capture contact angle variations. In (9), the strength factor $s_{xx}$ (with $xx$ denoting the phase pair, e.g., $A$, $B$, …, $n$), governs interfacial interactions and equilibrium configurations [25]. To illustrate, we consider a system of four phases where phase $1$ contains droplets of phases $2$, $3$, and $4$, initially non-contacting, as shown in detail in Figure 1(c). Interaction strengths are defined as $s_{11}=s_{22}=s_{33}=s_{44}$ and $s_{12}=s_{13}=s_{14}=s_{23}=s_{24}=s_{34}$. Under these conditions, the system evolves into an equilibrium configuration with a symmetric triple contact angle of $120^{\circ}$ between each pair of interfaces. Similarly, varying the interaction strengths should enable us to control the static contact angles on solid surfaces[15, 34]. In Figure 1(c), we present the contact angle variation between a fluid (1), a droplet (2) and a solid surface (3). To estimate the contact angle, we measure the droplet height $a_{d}$ and width $2b_{d}$, such that $\theta={\pi}/{2}-\arcsin({b_{d}^{2}-a_{d}^{2}}/{b_{d}^{2}+a_{d}^{2}})$ [25]. To obtain angles between $0^{o}$ and $90^{o}$ it is necessary that the interaction force between fluid-droplet be less than the interaction force between droplet-solid ($s_{13}<s_{23}$). In the other hand, for $90^{o}$ to $180^{o}$, it is necessary that the interaction force between fluid-droplet be greater than the interaction force between droplet-solid ($s_{13}>s_{23}$). We compare the measured angles between 0 and 90 plotted against the droplet-solid interaction force ($s_{23}$) normalized to its maximum value ($s_{23,max}$). In Figure 1(c), we show that the measured variation in the contact angle is consistenty with previously reported numerical results  [15, 34]. This static validation confirms that the implemented multiphase scheme reliably captures interfacial behavior and surface tension effects across fluid–fluid and fluid–solid boundaries.
@@ 3.2 Dynamic validation: Poiseuille and shear flow
In the following, we validate the multiphase model by simulating two benchmark cases: a Poiseuille flow in a channel with two fluid phases and the dynamics of a droplet suspended in a liquid domain. The first case allows us to assess the model’s ability to capture the velocity profile and interface stability under varying viscosity ratios, while the second case tests the model’s capability to simulate droplet deformation and breakup under shear flow conditions. First, we consider a Poiseuille flow in a channel containing two fluid phases, $A$ and $B$, initially distributed in the lower and upper half of the channel, respectively. We set a constant viscosity for the phase $A$ ($\eta_{A}=1$) and different the viscosities for the phase $B$, by varying viscosity ratio $\lambda=\eta_{B}/\eta_{A}$ in the range $\{1,2,5,1\}$. For this setting, the theoretical velocity profile follows [24]
$v(x)=\rho gL^{2}\begin{cases}-\frac{1}{2\mu_{B}}\left(\frac{y}{L}\right)^{2}+a\frac{1}{\mu_{B}}\frac{y}{L}&0<y<L/2,\\ \frac{1}{2\mu_{A}}\left(1-\left(\frac{y}{L}\right)^{2}\right)+a\frac{1}{\mu_{A}}\left(\frac{y}{L}-1\right)&L/2\leq y<L,\end{cases}$ (16)
where
$a=\left(\frac{3}{4\mu_{A}}+\frac{1}{4\mu_{B}}\right)\left(\frac{1}{\mu_{B}}+\frac{1}{\mu_{A}}\right)^{-1},$ (17)
@@
Next, we validate the model by simulating the dynamics of a droplet ($d$) of radius $r_{0}$, suspended in a continuous fluid ($f$) phase under shear flow conditions. The droplet is initially spherical and is subjected to a shear flow induced by a constant velocity gradient $\dot{\gamma}$. The capillary number $Ca=(\dot{\gamma}\ r_{0}\ \mu_{f})/\sigma$ quantify the relative effect of viscous forces to surface tension, where $r_{0}$ is the initial radius of the droplet, $\mu_{f}$ is the viscosity of the continuous phase.
In Figure 2(b), we present the measured Taylor deformation parameter $\mathcal{D}$ for different capillary numbers $Ca=\{0.05,0.1,0.2,0.3\}$. We consider a droplet of size $r_{0}=2.5$, within two walls moving at velocity $U$ and $-U$, respectively, in a square channel of size $L=6r_{0}$ The physical properties of the droplet and fluid are $\rho_{f}=\rho_{d}=1$, $\mu_{f}=\mu_{d}=5$ and $\sigma=2.5$. We vary $Ca$ by changing the walls velocity, leading to Reynolds number $Re=(\rho_{c}\ \dot{\gamma}\ r_{0}^{2})/\mu_{f}$ in the range $\{0.025,0.05,0.075,0.1\}$ always smaller than 1. In Figure 2(b), we can observe an excellent agreement of our numerical results with the theoretical predictions[41] for $Ca<0.4$, where $\mathcal{D}$ can be approximated as $D=(19\lambda+16)/(16\lambda+16)Ca$.
In Figure 2(c), we further explore the droplet dynamics as the capillary number increases, where the droplet undergoes breakup due to the unbalance between viscous forces and surface tension. Figure 2(c) shows the evolution from the undeformed state to the break-up of a droplet with $r_{0}=3$, $L=10r_{0}$, $\sigma=1$, $\mu_{f}=\mu_{d}=1$ and $\gamma=0.3$ corresponding to a $Ca=0.9$. The breakup occurs when the droplet’s aspect ratio exceeds a critical value, leading to the formation of smaller droplets. The breakup process is characterized by the formation of a neck at the droplet’s center, which eventually leads to the separation of the droplet into two smaller droplets. This behavior is consistent with previous studies on droplet dynamics in shear flows[41, 42, 43].
@@ 3.3 Thixotropic model validation
We validate our implementation of the thixotropic model considering a single-phase fluid in a simple shear flow within two parallel plates moving in opposite directions. For these simulations, we define a distance between plates $L=10$ with an initial interparticle distance $\Delta x=0.2$. The fluid properties are set as $f_{0}=1$, $\eta_{\infty}=15$, $\rho=1$, and $a=b=0.025$ (all within the adopted unit system), to fulfill the condition $\lambda_{0}>L^{2}\rho/\eta_{\infty}$ [62]. Following Rossi et al. [62], we track the evolution of the shear stress $\tau$ and the steady state viscosity $\eta$, from the measured tangential force on the walls $F_{x}$, such that $\tau={F_{x}}/{A}$ and $\eta(\dot{\gamma})={F_{x}}/{A\dot{\gamma}}$, where $A$ is the wall area.
In Figure 3, we summarize the results of the thixotropic model validation, evidencing the accuracy of the model in both steady and transient conditions. In Figure 3(a) and Figure 3(b), we present the variation in the measured steady state values of $\tau$ and $\eta$ (for various values of $\alpha=[1,2,4]$), as the shear rate $\dot{\gamma}$ increases, following an excellent agreement with the theoretical model Equation 8 (see Section 2). In Figure 3(c), we further present the transient behavior of $\eta(t)$ for a fluid with $\alpha=1$ under different shear rates. The results evidence the microstructural evolution of the fluid, where the viscosity decreases with time as the microstructure forms and grows at different rates. For the lowest values of shear rate, noticeable oscillations in the instantaneous force measurement lead to small fluctuation in the measured $\eta$. However, these numerical oscillations do not affect the overall trend of the viscosity evolution, which remains consistent with the expected behavior of a thixotropic fluid.
(a) Shear stress
(b) Steady shear viscosity
(c) Viscosity evolution for $\alpha=1$
Figure 3: Thixotropic model validation for a simple shear flow for $\alpha$ range of (1,2,4)). Model parameters are based on the validation process developed by [62]
@@ 4 Case Studies: Thixotropic Multiphase Dynamics from Phase Separation to Microfluidics
Following the full validation of the thixotropic multiphase SDPD methodology, we now apply the model to a series of exploratory case studies. These examples are designed to demonstrate the model’s versatility in capturing complex multiphase dynamics under different physical conditions. Specifically, we examine: (i) proteins liquid–liquid phase separation (LLPS), (ii) the behavior of droplet suspensions, (iii) droplet transport through a periodically constricted channel, and (iv) droplet coalescence within microfluidic devices. These scenarios showcase the model’s capability to handle both fundamental and applied problems involving thixotropic behavior and intricate interfacial interactions.
4.1 Liquid-Liquid Phase Separation (LLPS)
@@
Liquid–liquid phase separation is a wide-spread phenomena occurring in biological processes [44, 46, 91], such as the formation of protein droplets in cells to form membraneless organelles [92, 93], intrinsically disordered proteins phase separation in neurodegenerative diseases [88, 89, 90], to name a few. In LLPS the systems can evolve dynamically from homogeneously mixed protein solutions to heterogeneous multiphase-phase systems. The process is characterized by the formation of solute-rich droplets dispersed in a solute-poor continuous phase, where the microstructural properties of the dispersed phase can change over time due to different physical process (coaservation, gelation, cristalization, etc), leading to a thixotropic behavior. In this context, the thixotropic multiphase model can be used to describe the dynamics of proteins LLPS, where phase separation time scales $\lambda_{\sigma}$ and thixotropic time scales $\lambda_{0}$ are interwined.
@@
We investigate systems containing two type of particles, describing a disperse or ”protein phase” ($p$) and a continuous or ”solvent phase” ($s$), with volume fractions $\Phi_{p}$ and $\Phi_{s}=\Phi_{p}$, respectively. As seen in Figure 4, the system evolves from an initial homogeneous condition to a final steady state, where the protein phase separate into droplets. Conceptually, we can think of the protein phase as a collection of protein aggregates dispersed in a continuous solvent phase. The protein phase is characterized by a critical concentration of proteins that triggers the phase separation process, leading to the formation of protein-rich droplets. The solvent phase contains the majority of the solvent and is responsible for the transport and interaction of the protein droplets. Note that we do not model explicitly the protein molecules but rather consider the protein phase as a thixotropic fluid which viscosity changes as protein aggregates form.
For a given $\Phi_{d}$, we can expected that a characteristic time scale $\lambda_{\sigma}$ for the phase separation process lies between a coarsening time scale $R^{2}/D_{p}$ and the coalescence time scale $R\eta_{p}/\sigma_{o}$ ($R^{2}/D_{p}<\lambda_{\sigma}<R/Ca_{v}$), where $R$ is the mean radius of the droplets, $D_{p}\propto k_{B}T/(\eta h^{2})$ is the diffusion coefficient of the protein phase, and $Ca_{v}=\sigma_{o}/\eta_{p}$ is the capillary velocity[46] with $\eta_{p}$ being the viscosity of the protein phase, and $\sigma_{o}$ is the surface tension between the protein droplets and the solvent phase. The coarsening time scale $R^{2}/D_{p}$ represents the time it takes for droplets to grow by diffusion, while the coalescence time scale $R/Ca_{v}$ represents the time it takes for droplets to merge due to surface tension effects. The thixotropic time scale $\lambda_{0}$, governs the rate at which the protein fluid evolves into highly viscous aggregates. The viscosity of the protein phase is assumed to increase with time, with an initial microstructural scalar parameter $f_{0}=0$, as the protein aggregates form and grow. Thus, $\eta_{\infty}$ is the viscosity of the diluted protein phase in a liquid-like state and $\alpha$ determines the maximum viscosity of the protein phase, when $f=1$. To asses the interplay between the phase separation and thixotropic time scales, we investigate the effect of four different parameters on the LLPS process, named (i) protein phase volume fraction $\Phi_{d}$, (ii) capillary velocity $Ca_{v}=\sigma_{o}/\eta_{s}$, (iii) constitutive constant ${1}/{1+\alpha}$, and (iv) thermal energy $k_{B}T$.

### v2

@@ 3.1 Validation of the numerical methodology
We conduct a series of benchmark simulations to ensure the accuracy and robustness of our implementation. First addressing the correct description of the surface tension and contact angles in static and dynamic cases. Then, we validate the thixotropic model by simulating a transient viscosity case. The results of these simulations are compared with theoretical predictions and previous numerical results. Unless otherwise indicated, all the simulations presented use an initial interparticle distance $\Delta x=0.2$.
3.1.1 Static validation: Surface tension and contact angles
As is customary tartakovsky2005modeling , tartakovsky2006pore , tartakovsky2009lagrangian , kordilla2013smoothed , tartakovsky2016pairwise , lei2016smoothed , we initially validate the consistency of our numerical implementation to reproduce the surface tension ($\sigma$) of a spherical droplet in a fluid via the Young Laplace equation, $\sigma=R(P_{d}-P_{f})/2$, where $R$ is the droplet radius, and $P_{d}$ and $P_{f}$ are the pressure of the droplet and external fluid, respectively. For these simulations, we start with a square-shaped particle array of different sizes, that evolve in time to circular droplets of varying size. The square simulation box has a length $L/\Delta x=150$. In Equation 14, we use an input $\sigma_{0}=0.25$ and $\sigma_{0}=1$, with an equilibrium particle density $n_{eq}=25$, viscosity $\eta=1$ for both droplet and continuous phase and $k_{B}T=0$. These parameters lead to an interfacial tension time scale $\tau_{\sigma}=\{1,4\}$ for each of the $\sigma_{0}$ evaluated, considering a normalized radius droplet with unit length as the characteristic length, i.e., $R=1$. In Figure 1(a), we present the evolution of the pressure difference as a function of curvature ($1/R$). Consistent with the Young-Laplace equation, the slopes obtained from the linear fitting, $0.53$ and $2.08$ corresponds to $\sim 2\sigma$. For these surface tension output values, it can be calculated that $\tau_{\sigma}=\{0.96,3.78\}$ showing the consistency of the proposed methodology for reproducing this time scale.
We further validate the consistency of our method, measuring the deformation of an initially elliptical droplet as it retracts to a circular shape due to interfacial tension effects luciani1997interfacial , guido1999measurement , mo2000new , pan2015studies . The initial ellipse (at $t=0$) is defined by the semi-major axis $I_{A}$ and the semi-minor axis $I_{B}$, defining the so called Taylor deformation parameter taylor1934formation ${\mathcal{D}}_{0}=(I_{A}-I_{B})/(I_{A}+I_{B})$. The retraction process is governed by the balance between surface tension and viscous forces. The evolution of the of the Taylor parameter $\mathcal{D}$ can be described by luciani1997interfacial , guido1999measurement , mo2000new
$\mathcal{D}=\mathcal{D}_{0}\exp\left(-\frac{\sigma}{\eta R}\frac{40(\lambda+1)}{(2\lambda+3)(19\lambda+16)}t\right),$ (16)
@@ 3.1.1 Static validation: Surface tension and contact angles
where $R$ is the final circle radius, $\lambda=\eta_{d}/\eta$ the ratio between the viscosity of the disperse phase $\eta_{d}$ and the continuous phase $\eta$, and $t$ the stretching time. The system is confined in a square simulation box of side length $L/\Delta x=50$. Here, we use three different viscosity ratios $\lambda=[1,5,10]$, and use Principal Component Analysis (PCA) to compute the eigenvalues and eigenvectors of the droplets particles, providing the magnitude and orientation of the ellipse’s principal axes. In Figure 1(b), we compare our results for the time evolution of $\mathcal{D}$ with (16), evidencing a good agreement between the numerical and theoretical results.
@@
We now validate the model’s ability to consistently capture contact angle variations. In (11), the strength factor $s_{xx}$ (with $xx$ denoting the phase pair, e.g., $1$, 22, …, $n$), governs interfacial interactions and equilibrium configurations tartakovsky2016pairwise . To illustrate, we consider a system of four phases where phase $1$ contains droplets of phases $2$, 33, and $4$, initially non-contacting, as shown in detail in Figure 1(c). We consider a square box for the simulation, where the length of each side is $L/\Delta x=100$. Interaction strengths are defined as $s_{11}=s_{22}=s_{33}=s_{44}$ and $s_{12}=s_{13}=s_{14}=s_{23}=s_{24}=s_{34}$. Under these conditions, the system evolves into an equilibrium configuration with a symmetric triple contact angle of $120^{\circ}$ between each pair of interfaces. Similarly, varying the interaction strengths should enable us to control the static contact angles on solid surfaceskordilla2013smoothed , li2018smoothed . In Figure 1(c), we present the contact angle variation between a fluid (1), a droplet (2) and a solid surface (3). To estimate the contact angle, we measure the droplet height $a_{d}$ and width $2b_{d}$, such that $\theta={\pi}/{2}-\arcsin({b_{d}^{2}-a_{d}^{2}}/{b_{d}^{2}+a_{d}^{2}})$ tartakovsky2016pairwise . To obtain angles between $0^{o}$ and $90^{o}$ it is necessary that the interaction force between fluid-droplet be less than the interaction force between droplet-solid ($s_{13}<s_{23}$). In the other hand, for $90^{o}$ to $180^{o}$, it is necessary that the interaction force between fluid-droplet be greater than the interaction force between droplet-solid ($s_{13}>s_{23}$). We compare the measured angles between 0 and 90 plotted against the droplet-solid interaction force ($s_{23}$) normalized to its maximum value ($s_{23,max}$). In Figure 1(c), we show that the measured variation in the contact angle is consistenty with previously reported numerical results  kordilla2013smoothed , li2018smoothed . This static validation confirms that the implemented multiphase scheme reliably captures interfacial behavior and surface tension effects across fluid–fluid and fluid–solid boundaries.
@@ 3.2 Dynamic validation: Poiseuille and shear flow
In the following, we validate the multiphase model by simulating two benchmark cases: a Poiseuille flow in a channel with two fluid phases and the dynamics of a droplet suspended in a liquid domain. The first case allows us to assess the model’s ability to capture the velocity profile and interface stability under varying viscosity ratios, while the second case tests the model’s capability to simulate droplet deformation and breakup under shear flow conditions. First, we consider a Poiseuille flow in a channel containing two fluid phases, $1$ and $2$, initially distributed in the lower and upper half of the channel, respectively. We set a constant viscosity for the phase 11 ($\eta_{1}=1$) and different the viscosities for the phase $2$, by varying viscosity ratio $\lambda=\eta_{1}/\eta_{2}$ in the range $\{1,2,5,10\}$. The height of the channel is $L/\Delta x=50$. For this setting, the theoretical velocity profile follows tartakovsky2009lagrangian
$v(x)=\rho gL^{2}\begin{cases}-\frac{1}{2\eta_{2}}\left(\frac{y}{L}\right)^{2}+\bar{\eta}\frac{1}{\eta_{2}}\frac{y}{L}&0<y<L/2,\\ \frac{1}{2\eta_{1}}\left(1-\left(\frac{y}{L}\right)^{2}\right)+\bar{\eta}\frac{1}{\eta_{1}}\left(\frac{y}{L}-1\right)&L/2\leq y<L,\end{cases}$ (17)
where
$\bar{\eta}=\left(\frac{3}{4\eta_{1}}+\frac{1}{4\eta_{2}}\right)\left(\frac{1}{\eta_{2}}+\frac{1}{\eta_{1}}\right)^{-1},$ (18)
Next, we validate the model by simulating the dynamics of a droplet ($d$) of radius $R_{0}$, suspended in a continuous fluid ($f$) phase under shear flow conditions. The droplet is initially spherical and is subjected to a shear flow induced by a constant velocity gradient $\dot{\gamma}$. The system is confined in a square simulation box of side length $L/\Delta x=100$.
In Figure 2(b), we present the measured Taylor deformation parameter $\mathcal{D}$ for different capillary numbers Ca=[0.050.3]Ca=[0.05-0.3]. We consider a droplet of size $R_{0}/\Delta x=12.5$, within two walls moving at velocity $U$ and U-U, respectively, in a square channel of size $L=6R_{0}$ . The physical properties of the droplet and fluid are $\rho_{f}=\rho_{d}=1$, $\eta_{f}=\eta_{d}=5$ and $\sigma=2.5$. We vary $Ca$ by changing the walls velocity, leading to Reynolds number in the range $\{0.025,0.05,0.075,0.1\}$ always smaller than 1. In Figure 2(b), we can observe an excellent agreement of our numerical results with the theoretical predictions pan2014dissipative for $Ca<0.4$, where $\mathcal{D}$ can be approximated as $\mathcal{D}=(19\lambda+16)/(16\lambda+16)Ca$.
In Figure 2(c), we further explore the droplet dynamics as the capillary number increases, where the droplet undergoes breakup due to the unbalance between viscous forces and surface tension. Figure 2(c) shows the evolution from the undeformed state to the break-up of a droplet with $R_{0}/\Delta x=15$, $L=10R_{0}$, $\sigma=1$, $\eta_{f}=\eta_{d}=1$ and $\gamma=0.3$ corresponding to a $Ca=0.9$. The breakup occurs when the droplet’s aspect ratio exceeds a critical value, leading to the formation of smaller droplets. The breakup process is characterized by the formation of a neck at the droplet’s center, which eventually leads to the separation of the droplet into two smaller droplets. This behavior is consistent with previous studies on droplet dynamics in shear flows pan2014dissipative , wang2023dynamics , zhang2023shear .
@@ 3.3 Thixotropic model validation
(a) Steady shear viscosity
(b) Viscosity evolution for $\alpha=1$
Figure 3: Thixotropic model validation for a simple shear flow for $\alpha$ range of (1,2,4)). Model parameters are based on the validation process developed by rossi2022sph
We validate our implementation of the thixotropic model considering a single-phase fluid in a simple shear flow within two parallel plates moving in opposite directions. We analyze five particular cases for which we set a sonic time scale $\tau_{\text{c}}=0.5$, characteristic flow time scale $\tau_{\text{flow}}=\{10,3.33,2,1.43,1.11\}$ and thixotropic time $\lambda_{\text{thix}}=40$, following the hierarchy $\tau_{\text{c}}\ll\tau_{\text{flow}},\lambda_{\text{thix}}$. We define a distance between plates $L/\Delta x=50$ and a set of five velocities corresponding to $U=\{0.5,1.5,2.5,3.5,4.5\}$. The fluid properties are set as $f_{0}=1$, $\eta_{\infty}=15$, $\rho=1$, and $k_{f}=k_{d}=0.025$ (all within the adopted unit system), to fulfill the condition $\lambda_{\text{thix}}>L^{2}\rho/\eta_{\infty}$ rossi2022sph . These parameters result in a Reynolds number $Re=\{0.33,1,1.66,2.33,3\}$ and a Thixotropic number $Th=\{0.1,0.3,0.5,0.7,0.9\}$. Following Rossi et al. rossi2022sph , we track the evolution of the steady state viscosity $\eta$ from the measured tangential force on the walls $F_{x}$, such that $\eta(\dot{\gamma})={F_{x}}/{A_{w}\dot{\gamma}}$, where $A_{w}$ is the wall area.
In Figure 3, we summarize the results of the thixotropic model validation, evidencing the accuracy of the model in both steady and transient conditions. In Figure 3(a) we present the variation in the measured steady state values of $\eta/\eta_{\infty}$ (for various values of $\alpha=[1,2,4]$ in Equation 6), as the shear rate $\dot{\gamma}$ increases, following an excellent agreement with the theoretical model Equation 10 (see Section 2). In Figure 3(b), we further present the transient behavior of $\eta(t)$ for a fluid with $\alpha=1$ ($\eta_{\text{max}}/\eta_{\infty}=2$) under different shear rates. The results evidence the transient microstructure evolution, leading to larger microstructure destruction (lower viscosity) as the applied shear rate increase. For the lowest values of $\dot{\gamma}$, noticeable oscillations in the instantaneous force measurement lead to small fluctuation in the estimated $\eta$. However, these numerical oscillations do not affect the overall trend of the viscosity evolution, which remains consistent with the expected behavior of a thixotropic fluid.
This validation highlight the interplay between the characteristic time scales. The variation in the flow time scale $\tau_{\text{flow}}$ relative to the thixotropic relaxation time $\lambda_{\text{thix}}$ reveals that when $\tau_{\text{flow}}\ll\lambda_{\text{thix}}$, the fluid response is dominated by the slow restructuring of the microstructure, leading to significant transient viscosity effects. Conversely, for larger $\tau_{\text{flow}}$, the system would approach steady-state more quickly, and thixotropic effects would be less pronounced.
@@ 4 Case Studies: Thixotropic Multiphase Dynamics from Phase Separation to Microfluidics
We now demonstrate the model’s versatility in capturing complex multiphase dynamics under different physical conditions. Specifically, we examine: (i) proteins liquid–liquid phase separation (LLPS), (ii) the behavior of droplet suspensions, (iii) droplet transport through a periodically constricted channel, and (iv) droplet coalescence within microfluidic devices.
4.1 Liquid-Liquid Phase Separation (LLPS) and aging
@@
Liquid–liquid phase separation leading to droplets formation is a wide-spread phenomena occurring in biological processes alberti2017phase , alberti2019considerations , feng2019formation , such as the formation of organelles in cells volkov2024phase , courchaine2016droplet and coacervates in neurodegenerative diseases Gracia2022 , boyko2022tau , zbinden2020phase . The process is characterized by the formation of solute-rich droplets dispersed in a solute-poor continuous phase, where the microstructural properties of the dispersed phase can change over time due to different physical process (coacervation, gelation, cristalization, etc), leading to a thixotropic behavior. In this context, the thixotropic multiphase model can be used to describe the dynamics of proteins LLPS, where phase separation time scales $\lambda_{\sigma}$ and thixotropic time scales $\lambda_{\text{thix}}$ are interwined.
@@
We investigate systems containing two type of particles, describing a disperse or ”protein phase” ($p$) and a continuous or ”solvent phase” ($s$), with volume fractions $\Phi_{p}$ and $\Phi_{s}=1-\Phi_{p}$, respectively. As seen in Figure 4, the system evolves from an initial homogeneously mixed condition to a final steady state, where the protein phase separate into droplets. Conceptually, we can think of the protein phase as a collection of protein aggregates dispersed in a continuous solvent phase. The protein phase is characterized by a critical concentration of proteins that triggers the phase separation process, leading to the formation of protein-rich droplets. The solvent phase contains the majority of the solvent and is responsible for the transport and interaction of the protein droplets. Note that we do not model explicitly the protein molecules but rather consider the protein phase as a thixotropic fluid which viscosity changes as protein aggregates form.
@@
For a given $\Phi_{p}$, we can expected that a characteristic time scale $\lambda_{\sigma}$ for the phase separation process lies between a coarsening time scale $R^{2}/D_{p}$ and the ratio between the thixotropic and interfacial tension time scales $\lambda_{\text{thix}}/\tau_{\sigma}$. This is, $R^{2}/D_{p}<\lambda_{\sigma}<\lambda_{\text{thix}}/\tau_{\sigma}$, where $R$ is the mean radius of the droplets and $D_{p}\propto k_{B}T/(\eta_{s}h^{2})$ is the diffusion coefficient of the protein phase. The coarsening time scale represents the time it takes for droplets to grow by diffusion, while the ratio $\lambda_{\text{thix}}/\tau_{\sigma}$ represents the relation between the microstructural restructuring and capillary-driven coalescence. The thixotropic time scale $\lambda_{\text{thix}}$, governs the rate at which the protein fluid evolves into highly viscous aggregates. The viscosity of the protein phase is assumed to increase with time, with an initial microstructural scalar parameter $f_{0}=0$, as the protein aggregates form and grow. Thus, $\eta_{\infty}$ is the initial viscosity of the diluted protein phase.
(a) $\lambda_{\text{thix}}/\tau_{\sigma}$
(b) initial-to-final viscosity ${1}/{(1+\alpha})$
(c) Thermal fluctuations $k_{B}T$
Figure 5: KDE comparison for the different properties analyses including the PID value. Inside snapshot illustrate the final configuration for two different system, where droplets are depicted in different colors for clarity. We employ the properties values for (a) $\lambda_{\text{thix}}/\tau_{\sigma}$ with $\sigma=[0.1,0.5,1.0,2.0]$, $\alpha=1$ (i.e. $\eta_{\text{max}}/\eta_{\infty}=2$), $k_{B}T=0$, for (b) initial-to-final viscosity with $\sigma=2.0$, $\alpha=[1,4,9,99]$, $k_{B}T=0$ and for (c) Thermal fluctuations with $\sigma=2.0$, $\alpha=1$ (i.e. $\eta_{\text{max}}/\eta_{\infty}=2$), $k_{B}T=\{0,0.05,0.1\}$. For all simulations we use a fixed protein phase $\phi_{p}=0.25$, viscosity of the solvent phase as $\eta_{s}=1$, thixotropic time scale $\lambda_{thix}=1$ and $f_{0}=0$
To asses the interplay between the phase separation and thixotropic time scales, we investigate the effect of (i) variation of the ratio between the thixotropic characteristic time and interfacial tension time scale $\lambda_{\text{thix}}/\tau_{\sigma}=\{0.1,0.5,1,2\}$, (ii) ratio initial-to-final viscosity ${1}/({1+\alpha})=\{0.01,0.05,0.1,0.5\}$, and (iii) thermal fluctuations time scale $\tau_{t}=\{20,28\}$ obtained by the variation in the thermal energy $k_{B}T=\{0.05,0.1\}$ . To calculate the characteristic times, we use a normalized radius droplet with unit length as the characteristic length, i.e., $R=1$ for the case of $\tau_{\sigma}$, and the length of the box $L=10$ for the case of $\tau_{\text{t}}$. For simplicity, we study these parameters under a fixed protein phase concentration $\phi_{p}=0.25$. Note that since protein phase concentration $\phi_{p}$ can also affect the phase separation dynamics in a non-linear fashion. In B, we briefly illustrate the effect of $\phi_{p}$, where lower concentrations favor smaller droplets with narrower distributions, whereas larger concentrations lead to larger droplets or spinodal-type separation of the phases. We conduct LLPS simulations in square domains of size $L/\Delta x=50$, in quiescent conditions ($\dot{\gamma}=0$) such that no microstructural breaking is considered. We fix the viscosity of the solvent phase as $\eta_{s}=1$. All simulations reported in this subsection were conducted at a fixed thixotropic time scale $\lambda_{\text{thix}}=1$, thus, the variations observed correspond to changes in the relative magnitude of $\lambda_{\sigma}$ with respect to $\lambda_{\text{thix}}$.